OpenStudy (anonymous):

The aurora is caused when electrons and protons, moving in the earth's magnetic field of ≈5×10−5 T, collide with molecules of the atmosphere and cause them to glow. What is the radius of the circular orbit for (a) an electron with speed 1.0×106m/s and (b) a proton with speed 7.0×104m/s . What is re in meters and what is rp in meters?

OpenStudy (anonymous):

You can use this, qvB = (mv^2)/r

OpenStudy (anonymous):

remember that \[e=1.602x10^{-19}C\] \[q_p = +e\] \[q_e=-e\] And be sure to keep track of units. units on the magnetic field \[1 T = \frac{1 kg}{s^2A}\] and an Ampere can be written as \[1 A = \frac{1C}{s}\]